The demand and supply equations for a product are:Qd=300_6p and Qs-4+6p.

Determine the market equilibrium and draw graph.

Suppose that the government decides to impose a flat of Rs 1.70 per unit sold.Draw graph and explain.

Also calculate the total revenue earned by sellers before and after the text revenue raised by the government,changes in consumer and producers surplus and dead weight loss.


1
Expert's answer
2022-01-07T09:18:41-0500

Solution:

At equilibrium, Qd = Qs

300 – 6p = -4 + 6p

300 + 4 = 6p + 6p

304 = 12p

P = 25.3

Equilibrium price = 25.3

Substitute to derive equilibrium quantity:

Qd = 300 – 6p = 300 – 6(25.3) = 200 – 151.8 = 148.2

Equilibrium quantity = 148